What is the solubility of Fe(OH)3 ( iron(III) hydroxide) in...

a) ...pure water?


b) ...0.0005 M FeCl3?
Answer #1

& solubility equilibrium of Fe (OHC is Fe(OH) – Fostea + 3011 caq Ksp = [Fe3+] [A3 in pure water. Ksp for fe @H) = 6-3x10 38= 4.667 X10-36 Ksp = 6-3x10 38 = 0.0005 (35) = 0.0135 53 = 6.3 x10 38 466-67x100 0.0135 S = (4.667 X163673 = 1.6711 x 1612 so